OpenStudy (anonymous):

How do I simplify √(x-20)^2 + 5 √=square root and the 2 is outside of the square root

OpenStudy (anonymous):

√(x-20)^2=√(x-20)*√(x-20)

OpenStudy (anonymous):

\[[\sqrt{x-20}]^2+5=x-20+5=x-15\]